### v1

# From Instruction to Event: Sound-Triggered Mobile Manipulation

Abstract
Current mobile manipulation research predominantly follows an instruction-driven paradigm, where agents rely on predefined textual commands to execute tasks. However, this setting confines agents to a passive role, limiting their autonomy and ability to react to dynamic environmental events. To address these limitations, we introduce sound-triggered mobile manipulation, where agents must actively perceive and interact with sound-emitting objects without explicit action instructions. To support these tasks, we develop Habitat-Echo, a data platform that integrates acoustic rendering with physical interaction. We further propose a baseline comprising a high-level task planner and low-level policy models to complete these tasks. Extensive experiments show that the proposed baseline empowers agents to actively detect and respond to auditory events, eliminating the need for case-by-case instructions. Notably, in the challenging dual-source scenario, the agent successfully isolates the primary source from overlapping acoustic interference to execute the first interaction, and subsequently proceeds to manipulate the secondary object, verifying the robustness of the baseline.
Machine Learning, ICML
Figure 1: Motivation for our work. (Upper) An example of sound-triggered mobile manipulation, the trigger signal is the sound of the event rather than instructions from humans. (Left) Instruction-based mobile manipulation needs humans to analyze the trigger signal and manually give the instruction for the robot model, which is passive. (Right) In sound-triggered mobile manipulation, the robot model automatically receives sound signals and actively outputs executable actions.
1 Introduction
Recent advances in Embodied AI have witnessed remarkable progress in mobile manipulation, enabling robots to navigate unstructured environments and interact with physical objects (Song et al., 2025a; Bai et al., 2025; Yu et al., 2025). Such synergy between mobility and interaction empowers agents to perform long-horizon household tasks such as object rearrangement (Ji et al., 2025; Huang et al., 2025; Escoriza et al., 2025) and multi-stage navigation (Song et al., 2025b; Lin et al., 2025), bringing us closer to general-purpose home robotic assistants.
Current paradigm predominantly studies mobile manipulation under an instruction-conditioned setting, where the task goal and intended interaction are explicitly specified by a natural language instruction (Wu et al., 2025; Shah et al., 2025; Yan et al., 2025; Zhang et al., 2025). While effective for user-specified chores, this paradigm is ill-suited for the stochastic nature of daily life. For example, events such as a boiling kettle or a ringing doorbell typically happen unexpectedly and require timely intervention, where the latency inherent in human-in-the-loop interaction renders timely intervention impossible. This uncertainty is further compounded when multiple events occur concurrently or in rapid succession, i.e., waiting for a phone call from friends while expecting a doorstep delivery from a courier simultaneously. The indeterminate sequence of such events prevents the user from formulating a valid instruction plan in advance.
This limitation underscores the discrepancy between static, predefined instructions and dynamic, sound-triggered scenarios (See Figure 1). Although recent efforts have attempted to generalize instructions ranging from concrete commands (Yenamandra et al., 2023) to abstract goals (Wu et al., 2023) and ambiguous queries (Ichter et al., 2022; Li et al., 2025), addressing these stochastic events remains an open challenge. In this context, we find that sound signals play a pivotal role. For instance, if the phone rings prior to a knock at the door, the robot should prioritize picking up the phone before retrieving the delivery. In such cases, robotic actions are modulated by stochastic sounds rather than predefined text instructions. Specifically, sounds serve a threefold function: (1) temporal trigger indicating the start of a stochastic event; (2) semantic cue guiding the robot’s action selection; and (3) spatial cue revealing the event’s location. Based on these insights, we introduce sound-triggered mobile manipulation, comprising SonicStow, SonicInteract, and Bi-Sonic Manipulation, where acoustic signals replace conventional textual instructions as the trigger signal. SonicStow and SonicInteract present unique challenges beyond traditional visual manipulation: (1) Acoustic semantic understanding, which requires the robot to interpret sounds to determine target locations and actions. (2) Sound-conditioned interaction, which demands the execution of articulated interactions solely triggered by audio. The Bi-Sonic Manipulation task further elevates the complexity by introducing dual sound sources, which demands the agent to prioritize the primary source amidst acoustic interference and subsequently locate and interact with the secondary sound source.
To validate our proposed tasks, we introduce Habitat-Echo, a simulation platform that seamlessly integrates acoustic rendering with interactive physical environments. Built upon the visual and physical capabilities of Habitat (Szot et al., 2021; Puig et al., 2024), Habitat-Echo addresses the absence of sound signals in prior simulators. Specifically, we leverage room impulse responses to convolve source sound with the environment’s geometry, thereby ensuring physically accurate sound propagation and spatialization from the source to the receiver. Leveraging this platform, we construct a benchmark tailored for sound-triggered mobile manipulation. The dataset encompasses many interactive assets, categorized into: (1) static silent rigid objects, (2) sound-emitting rigid objects, and (3) sound-emitting articulated objects. We utilize this dataset to train policy models and further introduce an Omni-modal Large Language Model (Omni-LLM) as a high-level task planner to understand acoustic events and decompose complex manipulation goals. The main contributions of this work are summarized as follows:
  • Novel Sound-Triggered Tasks. Distinct from traditional instruction-following paradigms, we formulate three novel manipulation tasks triggered by stochastic acoustic signals. In these tasks, robotic actions are autonomously triggered by stochastic sound events rather than static predefined instructions, better reflecting the unpredictable and dynamic nature of real-world household scenarios.
  • New Audio-Visual-Physical Simulation Platform. We introduce Habitat-Echo, a platform that integrates acoustic and visual rendering with interactive physical environments. By incorporating sound-emitting objects, Habitat-Echo facilitates large-scale episode generation and closed-loop online interaction, establishing a robust foundation for future research in embodied audio-visual learning.
  • Hierarchical Baseline and Validation. We develop a hierarchical framework that is composed of a high-level task planner to interpret acoustic cues and several low-level executors for precise control. We leverage an Omni-LLM as the task planner and policy models as executors. The baseline framework achieves promising performance on the proposed benchmark.
Figure 2: Illustration of the proposed tasks. (a) SonicStow requires the agent to interact with the sound source (a rigid object) via Navigate, Pick, and Place skills. (b) SonicInteract requires the agent to interact with the sound source (an articulated object) via Navigate, Open Door, and Close Sink skills. (c) Bi-Sonic Manipulation requires the agent to interact with the sound source (two objects sequentially) via Navigate, Pick, Place, Open Door, and Close Sink skills.
2 Related Work
Sound Navigation. Sound Navigation, proposed by (Chen et al., 2020), requires an agent to navigate to an audio source without explicit goal coordinates. The primary challenges involve understanding physical spaces, localizing sound emitters, and avoiding obstacles. Existing methods can be broadly categorized into two streams: real-world application and versatile training. For real-world application, researchers aim to bridge the simulation-to-reality gap by introducing complex environmental settings. (Chen et al., 2021) assign semantic meanings to audio sources and introduce dynamic sound emission, replacing the continuous, generic signals used in prior work. (Younes et al., 2023) further extends the setting from static to moving sound sources. To simulate complex acoustic environments, (Yu et al., 2022) introduces an adversarial sound attacker that generates distracting noises with varying categories and volumes. For versatile training, methods leverage cross-modal synergies to enhance performance. VXN (Wang et al., 2022) unifies diverse navigation tasks, using sound navigation to facilitate the training of visual-language navigation. Recently, (Kondoh and Kanezaki, 2025) incorporates language descriptions of past and future actions as an auxiliary task to enhance policy learning. However, these works mainly focus on navigation capabilities, neglecting physical interactions with objects within the environment.
Mobile Manipulation. Mobile manipulation extends the robot’s workspace from confined tabletops to large-scale environments, requiring the synergy of mobility and interaction. Existing methods fall into two primary streams. The first stream focuses on object rearrangement in structured environments. (Szot et al., 2021) and (Shukla et al., 2024) develop high-fidelity simulators to support complex physical training and evaluation. Building on such platforms, (Gu et al., 2023) proposes a hierarchical framework sequencing navigation and interaction skills, while (Kant et al., 2022) integrates commonsense reasoning to infer logical object placements. The second stream targets open-vocabulary manipulation via VLMs. (Zitkovich et al., 2023) transfers web-scale knowledge to robotic control for generalization. To handle unseen objects, (Yenamandra et al., 2023) establishes a rigorous benchmark, and (Liu et al., 2024) achieves zero-shot performance by combining VLMs with navigation primitives. However, these methods remain predominantly instruction-driven. Relying on predefined textual commands limits the agent’s ability to automatically process and actively respond to dynamic events in the environment.
Task Planning for Robotic Manipulation. Task planning bridges high-level reasoning with low-level execution. Early approaches (Kaelbling and Lozano-Pérez, 2011; Toussaint, 2015; Srivastava et al., 2014) primarily relied on predefined symbolic representations for continuous control. Recent research leverages the extensive world knowledge within LLMs and VLMs for planning, broadly categorized into two streams The first stream focuses on task decomposition. Wake et al. (Wake et al., 2024) utilize GPT-4V (OpenAI, 2023) to ground spatial-temporal information within the demonstration video and analyze affordances for plan generation. Similarly, Kachaev et al. (Kachaev et al., 2025) employ VLMs to decompose instructions into subtask sequences. RoboBrain (Ji et al., 2025) unifies planning with affordance perception and trajectory prediction within a single model, enhancing manipulation capabilities. The second stream emphasizes feedback-based refinement. Zhi et al. (Zhi et al., 2025) optimize code-based plans using execution feedback, and Feng et al. (Feng et al., 2025) improve plans by anticipating future states. Similarly, Wang et al. (Wang et al., 2025) employ VLMs to solve PDDL problems constructed from current observations. However, these works primarily rely on visual and textual inputs, neglecting the importance of audio signals.
3 Task Definition and Skill
3.1 Task Definition
We formally define three novel sound-triggered tasks: SonicStow, SonicInteract, and Bi-Sonic Manipulation. SonicStow and SonicInteract are mobile manipulation tasks with a single sound source. Bi-Sonic Manipulation, a scenario with dual sound sources. See Figure 2 for more details.
SonicStow. This task simulates a scenario where an agent actively perceives an acoustic signal, localizes its source, and performs Pick and Place operations on household objects. Distinct from standard instruction-following paradigms that rely on predefined instructions, the agent must autonomously identify the target and conduct interaction solely based on the stochastic audio event. In each episode, the agent is initialized with a random pose and start location, while a single sound source, randomly selected from three objects (Phone, Alarm, Furby) within the scene (see Figure 2(a)). To introduce visual clutter, two distractor objects are randomly sampled from the YCB dataset for each episode. All objects are placed on elevated receptacles (e.g., existing furniture or inserted tables) instead of floor.
SonicInteract. This task entails localizing an acoustic event and performing manipulation on articulated objects. Different from picking and placing objects in SonicStow, this task requires the agent to precisely control the end-effector and manipulate articulated mechanisms, i.e., opening a door or turning off a faucet in the sink (see Figure 2(b)). In each episode, the agent is initialized with a random pose and start location. The target single-sound source is randomly sampled from two categories: a doorbell or a sink with running water. Object placement follows category-specific constraints: doorbells are fixed to the existing doors within the scene, whereas sinks are placed on the floor with arbitrary orientations and locations.
Bi-Sonic Manipulation. The Bi-Sonic Manipulation task integrates the challenges of SonicStow and SonicInteract, requiring the agent to manipulate both rigid and articulated objects (see Figure 2(c)). Additionally, this task introduces three more challenges. The first challenge of this task is the presence of two simultaneous sound sources, which creates a complex acoustic field featuring signal dynamics and superposition. Unlike single-source tasks, the agent must distinguish and interact with both sources sequentially based on a given priority (i.e., the main source followed by the secondary source). The categories and spatial locations of these sources are randomized. The second challenge of this task is two-time interactions. After interacting with the first object, the sound of the first object stops, and there is only one sound source in the scene. The agent has to execute skills and interact with the second object. Furthermore, to enhance visual complexity, we still introduce unrelated objects (Calli et al., 2015) to serve as visual clutter.
3.2 Skill
Similar to previous work (Gu et al., 2023; Shukla et al., 2024; Wang et al., 2025), the proposed tasks are decomposed into different subtasks. Each subtask is solved by an individual skill. A task is considered completed successfully if and only if all constituent skills are successfully executed in sequence. There are five skills in the proposed three tasks.
Navigate. The action space consists of four actions: MoveForward, TurnLeft, TurnRight, and Stop. Specifically, MoveForward translates the agent by 0.5m along its local forward axis, while rotational actions (TurnLeft, TurnRight) execute $90^{\circ}$ in-place rotations to the left and right, respectively. The Stop action terminates the episode without changing the agent’s pose and location. An episode is considered successful if and only if the agent executes Stop while positioned within a region of 0.5m radius of the goal.
Pick. The agent operates in an 8-DoF continuous action space. The first seven dimensions control the arm’s joint angles, driving the end-effector to the target pose via the kinematic chain (Abdelmaksoud et al., 2024), while the final dimension modulates the gripper state. A Pick skill is considered successful with two conditions: (1) Grasping: The gripper action is positive while an object is within a 15cm snapping threshold of the end-effector, triggering a stable grasp. (2) Retraction: The agent effectively holds the object and returns the arm to a predefined resting pose, ensuring a smooth transition to subsequent subtasks.
Place. The agent operates in an 8-DoF action space identical to that of the Pick skill. The success conditions for the Place skill are twofold: (1) Releasement: The gripper opens to release the object, ensuring it settles within 0.15cm of the target position. (2) Retraction: Upon release, the arm reverts to the default resting pose, preparing for the next subtask.
Open Door. Operating within the 8-DoF action space, the agent’s end-effector attaches to the handle of the door and drives it along its own functional trajectory. The skill is considered successful with two conditions: (1) Opening: While connected with the door handle, the agent must drive the door’s joint angle to a specified target threshold. (2) Retraction: The agent disengages from the door handle and returns the arm to the resting pose.
Close Sink. Similar to the Open Door skill, the agent utilizes the 8-DoF action space to manipulate the sink’s rotary mechanism. The end-effector must grasp the faucet handle and apply torque to rotate it along its functional axis. Success is determined by two conditions: (1) Closure: While maintaining grasp, the agent must rotate the handle to the zero-degree threshold within a tolerance of 0.2 rad. (2) Retraction: Upon completing the rotation, the agent releases the handle and reverts the arm to the resting pose.
Table 1: (a) Comparison of different simulators. Our Habitat-Echo supports sound rendering and physical interaction at the same time. (b) Distribution of audio assets across different categories in Habitat-Echo. We ensure a diverse split of sound instances for training and test sets.
Simulator Sound Rendering Physical Interaction
Habitat (Szot et al., 2021)
AI2-THOR (Kolve et al., 2017)
Isaac (Makoviychuk et al., 2021)
SoundSpaces (Chen et al., 2020)
Habitat-Echo (Ours)
Split Alarm Furby Phone Sink Doorbell Total
Train 11 30 19 6 11 77
Test 5 14 9 3 5 36
Total 16 44 28 9 16 113
4 Simulator and Benchmark
4.1 Simulator: Habitat-Echo
Existing simulators (Kolve et al., 2017; Makoviychuk et al., 2021; Szot et al., 2021) prioritize physical interaction with visual observations, while others (Chen et al., 2020, 2022) incorporate audio-visual rendering specifically for navigation tasks (see Table 1(a)). There remains an absence of simulators that integrate sound rendering with physical interaction. We address this limitation by introducing Habitat-Echo, an extension of Habitat (Szot et al., 2021) that enables sound rendering alongside physical interaction.
To enable high-fidelity acoustic simulation within the environment, we leverage precomputed Room Impulse Responses (RIRs) in SoundSpaces (Chen et al., 2020). An RIR serves as an accurate transfer function characterizing the multipath propagation between a source and a receiver (Kuttruff, 2016). RIRs provide high-fidelity acoustic features by incorporating two critical physical factors: (1) Geometry-aware propagation: Generated via bidirectional path tracing, RIRs implicitly capture intricate wave phenomena, including direct paths, specular reflections, and reverberation. In this way, rendered sounds are consistent with the room’s 3D geometry. (2) Material-dependent scattering: The RIR computation accounts for acoustic material properties by mapping semantic annotations to frequency-dependent absorption and transmission coefficients. This enables the simulator to reflect distinct acoustic interactions with diverse surfaces. With the help of RIRs, Habitat-Echo synthesizes the auditory observation by performing time-domain convolution between the source waveform and the corresponding RIR on the fly. Furthermore, exploiting the property of the acoustic field, our simulator supports simultaneous multi-source rendering by superimposing the convolved signals from distinct sources, which is essential for the Bi-Sonic Manipulation task. The final output is rendered as binaural audio, providing the agent with realistic spatial cues analogous to human auditory perception.
4.2 Benchmark
Sounding Objects and Audio Assets. In addition to standard YCB objects (Calli et al., 2015), we introduce five more categories of interactive sounding objects. For SonicStow, we introduce graspable rigid objects, including phone, alarm, and furby. These objects can be manipulated via Pick and Place skills. For SonicInteract, we incorporate articulated objects, including the sink and door. Unlike rigid items, these objects possess kinematic joints and are constrained to move along fixed functional axes. For the sink, we incorporate a dynamic water flow that is physically coupled to the faucet handle, moving synchronously with the handle’s rotation. To foster robust acoustic generalization, we also collect a diverse library of real-world sounds, implementing a strict one-to-many mapping where each object category is associated with multiple unique sound instances (see Table 1(b)). This design forces the agent to learn semantically meaningful audio-visual associations rather than overfitting to specific frequency patterns or waveforms, thereby ensuring high acoustic diversity of the benchmark.
Episode Generation. We employ a procedural randomization process to generate episodes for training and testing. In each episode, the sound for the target object is determined by hierarchically sampling a sound category and then a specific audio instance. Visual complexity is introduced by instantiating random silent objects as distractors alongside the target object. The agent is initialized at a random position, and target objects (excluding two doorbells in the scene) are also placed at random positions within the scene. We also validate the solvability of every episode by ensuring a traversable path connects the start and goal, and we strictly verify that interactive objects have accessible frontal regions for valid manipulation, especially for sinks and doors. An episode is considered successful if and only if the required skills are successfully executed in sequence.
5 Baseline
Figure 3: Overview of the proposed baseline. The sound-triggered task planner processes initial audio-visual observations to reason and generate a high-level skill chain from the skill library. Guided by this chain, specialized policy models are sequentially activated to generate low-level actions and interact with the environment.
The proposed baseline comprises two core components, i.e., Sound-triggered Task Planner and Policy Models (see  Figure 3). Sound-triggered task planner acts as the high-level reasoner, analyzing initial observations to predict a logical skill chain. Following this chain, the policy models serve as the low-level executors, sequentially translating each assigned skill into concrete actions to complete the corresponding subtasks.
5.1 Sound-triggered Task Planner
Motivated by MCP (Anthropic, 2024), we propose a sound-triggered task planner that operates as the high-level reasoning module. This module is responsible for translating environmental cues into a plan of skills, i.e., skill chain (see  Figure 3). The input for the task planner is twofold, comprising initial sensory observations and a skill library. Regarding observations, the planner utilizes the sound to reason about the target’s category and relative orientation, while the head RGB provides an initial egocentric view to perceive the scene layout partially. Crucially, to prevent information leakage, we restrict the planner’s input to strictly align with the observation space of the Navigate skill, ensuring no additional information is accessed. Complementing these observations, the skill library is provided as a textual prompt for the task planner. The library defines the semantic functionality of all available skills to help the reasoning of the planner. By combining these multi-modal observations with the skill descriptions, the task planner outputs an ordered chain of skills. This chain acts as a high-level directive, guiding the sequential activation of the following policy models to complete the task.
5.2 Policy Model
Policy models serve as low-level executors, translating the skill chain from the task planner into concrete actions and sequentially interacting with the environment. Policy models in the proposed baseline share a similar architecture comprising modality encoders, a temporal module, and actor-critic heads. The temporal module and actor-critic modules maintain a consistent structure following prior works (Chen et al., 2020; Gu et al., 2023), and the modality encoders are tailored to the specific observation space of different skills. For the Navigate skill, the observation space includes audio and head RGB, and the modality encoders are designed based on task complexity. In SonicStow and SonicInteract, the audio observations are converted into spectrograms and processed by a vanilla CNN encoder (Chen et al., 2020). In Bi-Sonic Manipulation, audio observations are transformed into Mel-spectrograms, and the encoder is LanguageBind (Zhu et al., 2024). An additional text encoder is introduced to specify the target sound source. For manipulation skills, including Pick, Place, Open Door, and Close Sink, the observation spaces consist of head and arm depth, which are processed by a standard CNN encoder (Chen et al., 2020).
Table 2: Results on the Habitat-Echo benchmark. We report the success rate ($\%$) of task planning, individual skills, and overall performance across different tasks. The higher the better. (a) Evaluation of SonicStow. (b) Evaluation of SonicInteract. (c) Evaluation of Bi-Sonic Manipulation. Success rates are reported in the format “First Sound Source / Second Sound Source”.
Baseline Task Planning Navigate Pick Place Overall
Individual N/A 57.66 88.73 70.27 N/A
Oracle 100.0 57.66 86.72 81.98 40.99
Qwen2.5-Omni-3B 12.16 55.56 93.33 71.43 4.50
Qwen3-Omni-30B 68.92 55.56 87.06 72.97 24.33
Qwen2.5-Omni-7B 78.38 51.72 85.56 79.22 27.48
Method Task Planning Navigate Open Door Close Sink Overall
Individual N/A 45.50 62.16 77.67 N/A
Oracle 100.00 45.50 50.00 72.63 27.90
Qwen2.5-Omni-3B 44.14 3.00 50.00 0.00 0.33
Qwen3-Omni-30B 59.01 44.70 33.33 73.20 14.05
Qwen2.5-Omni-7B 71.17 44.63 50.00 74.16 19.72
Method Task Planning Navigate Pick Place Open Door Close Sink Overall
Individual N/A 9.58 / 56.61 85.16 / 17.39 32.48 / 6.00 67.12 / 58.66 42.87 / 38.00 N/A
Oracle 100.00 9.58 / 46.67 92.31 / 0.00 25.00 / 0.00 72.73 / 0.00 40.00 / 100.00 4.23 / 0.56
Qwen2.5-Omni-3B 14.08 8.00 / 0.00 100.00 / 0.00 0.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.56 / 0.00
Qwen2.5-Omni-7B 17.18 14.75 / 100.00 100.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.00 / 100.00 0.56 / 0.56
Qwen3-Omni-30B 51.83 11.96 / 0.00 100.00 / 0.00 0.00 / 0.00 66.67 / 0.00 37.50 / 0.00 2.53 / 0.00
Figure 4: Qualitative Visualization of Task Execution. We present the execution trajectories for (a) SonicStow, (b) SonicInteract, and (c) Bi-Sonic Manipulation. For each task, the left panel illustrates the top-down view of the environment, highlighting the agent’s trajectories and the location of sound sources. The right panel displays the corresponding sequence of third-person keyframes during the execution. Note that Bi-Sonic Manipulation involves navigating to and interacting with two distinct sound sources sequentially.
6 Experiment
6.1 Implementation
Implementation Detail. We optimize all policy models using the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) on a single NVIDIA RTX 4090 GPU. The training process spans $5\times 10^{5}$ environment steps. Structurally, the visual and audio encoders remain frozen to ensure efficiency, and the Actor and Critic networks are Multi-Layer Perceptrons (MLPs) with a hidden dimension of 512. We utilize the Adam optimizer (Kingma, 2014) with an initial learning rate of $3\times 10^{-4}$ enabled with linear decay, a momentum of $1\times 10^{-5}$, and a maximum gradient norm of 0.5. Regarding PPO hyperparameters, we set the discount factor of 0.99, GAE parameter of 0.95, and clipping range of 0.2. The update phase involves 2 epochs with 2 mini-batches, using a rollout length of 128 steps. The value loss coefficient is set to 0.5, and the entropy coefficient is 0.1. For each task, the dataset is split into 660 instances for training. The testing dataset comprises 222 instances for both SonicStow and SonicInteract, and 355 instances for Bi-Sonic Manipulation.
Baseline. We train low-level policy models individually and then compose them with varying task planners, implementing three types of baselines. First, the Individual baseline evaluates each skill in isolation. This baseline bypasses the task planner to measure each skill without the influence of cascading errors. Second, the Oracle baseline employs a privileged task planner that provides the ground-truth skill chains, serving as a performance upper bound for the task planning. Finally, we integrate mainstream Omni-LLMs as task planners to evaluate audio-visual reasoning capabilities, including Qwen2.5-Omni (3B, 7B) (Xu et al., 2025) and Qwen3-Omni-30B-Instruct (Yang et al., 2025). These task planners are all training-free. We refer to these baselines directly by their own names.
6.2 Result
Quantitative Result of SonicStow. We present the evaluation results on the SonicStow task in Table 2(a). The Oracle baseline, which utilizes ground-truth planning, establishes the performance upper bound with an overall success rate of 40.99%. The Individual baseline reflects the standalone success rate of each policy model (e.g., 88.73% for Pick), decoupling low-level execution and high-level planning. Regarding task planning, Qwen2.5-Omni-7B achieves the highest accuracy of 78.38%, surpassing both the smaller Qwen2.5-Omni-3B and the larger Qwen3-Omni-30B. Driven by this, Qwen2.5-Omni-7B generates the highest number of correct skill chains, thereby attaining the best overall execution performance of 27.48% among all task planners.
Quantitative Result of SonicInteract. The evaluation results on the SonicInteract are shown in Table 2(b). With ground-truth task planning, the Oracle baseline achieves a 27.90% overall success rate, which is the theoretical upper bound of the policy models. Among the evaluated task planners, Qwen2.5-Omni-7B also emerges as the most effective one. It attains a planning rate of 71.17%, outperforming both Qwen2.5-Omni-3B and Qwen3-Omni-30B. The success rate of planning directly translates to execution success, and Qwen2.5-Omni-7B achieves the top overall performance of 19.72%. Conversely, we observe unsatisfactory cases in Qwen2.5-Omni-3B during the Close Sink subtask. Instead of responding to the sound of running water with the appropriate skill, Qwen2.5-Omni-3B consistently hallucinates a Pick and Place sequence, resulting in a near-zero success rate for this specific task.
Quantitative Result of Bi-Sonic Manipulation. Table 2(c) presents the results for the Bi-Sonic Manipulation task. In this dual-source environment, Qwen3-Omni-30B achieves the highest task planning accuracy, showing superior robustness against acoustic distraction. In contrast, Qwen2.5-Omni-7B fails to maintain its dominance in the task planning, receiving performance comparable to Qwen2.5-Omni-3B. Regarding the overall success rate, the Oracle and Qwen3-Omni-30B secure the top two results for the first sound source interaction. For the secondary source interaction, only the Oracle and Qwen2.5-Omni-7B manage to achieve 2 successful episodes. This near-zero success rate highlights the significant challenges posed by the Bi-Sonic Manipulation task (see Section. 3.1).
Table 3: Ablation Study on Input Modalities for Different Policy Models. We investigate optimal input combinations across three representative policy models for different skills, including Navigate, Pick, and Open Door. SR denotes success rate. (a) Policy model for Navigate. We evaluate different modality encoders including DINOv2 (Oquab et al., 2024) and CNN encoder (Chen et al., 2020) for visual inputs. Results indicate that combining Head RGB and Head Depth yields the highest success rate. (b) Policy model for Pick. Using a same encoder, the combination of Head Depth and Arm Depth outperforms single-view inputs. (c) Policy model for Open Door. Similarly, integrating Head Depth and Arm Depth is also proven effective for achieving optimal performance in the Open Door skill.
Model Input Modality Modality Encoder SR ($\uparrow$)
Head RGB Head Depth DINOv2 CNN
I 57.66
II 58.01
III 51.80
IV 61.71
Model Input Modality SR ($\uparrow$)
Head Depth Arm Depth
I 57.20
II 66.66
III 88.73
Model Input Modality SR ($\uparrow$)
Head Depth Arm Depth
I 46.39
II 47.29
III 62.16
Qualitative Result. Figure 4 presents qualitative visualizations of the agent’s execution across the three proposed tasks. In each subfigure, the left panel depicts the top-down trajectory and sound source locations, while the right panel displays sequential third-person keyframes of the interaction. Figure 4 (a) shows the SonicStow task. Initially, the target object (an alarm) is located outside the agent’s field of view. Relying on sound signals for localization, the agent successfully navigates to the target and executes the Pick and Place skills. Figure 4 (b) illustrates the SonicInteract task. The agent accurately navigates to the sound-emitting door and performs the Open Door skill. Figure 4 (c) showcases the Bi-Sonic Manipulation task, which involves two simultaneous sound sources: a primary source (yellow) and a secondary source (green). Guided by the task planner, the agent prioritizes the primary source despite the acoustic interference. It first navigates to and opens the door. Once the primary source is interacted, the agent isolates the remaining secondary signal, navigates to the sink, and executes the Close Sink skill. More visualizations can be seen in supp.
6.3 Ablation Study and Further Discussion
Effect of Input Modality for the Navigate Skill. We evaluate the impact of different visual modalities, including head RGB, head depth, and their corresponding encoders for the Navigate skill (see Table 3(a)). First, the model relying solely on head depth with a CNN encoder yields the lowest performance. This underperformance arises because depth observations lack semantic texture, making it difficult for the model to learn meaningful associations between visual cues in depth observations and the semantic information contained in audio observations. Second, models using head RGB achieve better results, with the trainable CNN encoder performing slightly better than the frozen DINOv2 due to its training process that helps the model to adapt to the indoor environment. Third, integrating both head RGB and head depth yields the best overall performance. This combination proves effective because the modalities are complementary. Head RGB provides semantic context to align with audio sources, while depth offers precise geometric distance information. Despite the performance of the dual-modality method, we select head RGB with the DINOv2 encoder as our default configuration to maintain an optimal balance between performance and training efficiency.
Effect of Input Modality for the Pick Skill. We investigate the impact of depth modalities on the Pick skill in Table 3(b), comparing the efficacy of head depth and arm depth. Using head depth alone yields the lowest success rate of 57.20%, primarily because the target object often falls outside the head camera’s optimal viewing range during manipulation. In contrast, the model relying solely on arm depth achieves a higher success rate of 66.66%, as the arm depth can directly observe the target when aligned. We observe that integrating both depth sources results in the best performance (88.73%). The two views prove complementary: the arm depth provides precise object-to-gripper distance, while the head depth offers global context to assist the arm in returning to the resting pose.
Effect of Input Modality for the Open Door Skill. For the Open Door skill, we validate the impact of different depth inputs, specifically analyzing head depth and arm depth (see Table 3(c)). Models relying on a single depth source achieve comparable performance, with the head depth only model (Model II) slightly outperforming the arm depth only model (Model I). We attribute this margin to the fact that head depth provides a global view of the door’s structure, whereas the arm camera often suffers from initial misalignment with the door’s specific affordance (i.e., the handle). Crucially, combining both modalities (Model III) results in a dramatic performance boost to 62.16%. This improvement stems from the complementary nature of the inputs: the head depth offers the global context necessary to locate the door frame, which in turn allows the agent to leverage the arm depth for precisely perceiving the target affordance for accurate manipulation.
7 Conclusion
In this work, we introduce sound-triggered mobile manipulation, a novel paradigm that shifts agents from passive instruction-following to active, sound-driven interaction. To support this paradigm, we develop Habitat-Echo, a data platform integrating audio rendering with physical interaction, and establish a hierarchical baseline comprising a high-level task planner and low-level policies. Extensive experiments validate that the proposed baseline enables agents to actively localize and manipulate sound-emitting objects. Notably, in challenging dual-source scenarios, the baseline demonstrates the capability to interact with the primary source amidst acoustic interference from a secondary signal, and successfully proceed to manipulate the subsequent target.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Qualitative Result
We present qualitative visualization results for the four manipulation skills in Figure 5. Figure 5 (a) and (b) illustrate the Pick skill. In the success case (a), the agent grasps the object and correctly returns the arm to the resting position, while in the failure case (b), the arm fails to retract to the standard pose despite holding the object. The Place skill is depicted in Figure 5 (c) and (d). In Figure 5 (c), the object is stably placed on the table, whereas in (d), the object is released with an unstable pose and falls off the surface of the small table. Figure 5 (e) and (f) visualize the Open Door skill. In the success scenario (e), the arm pulls the door open and resets to the resting position, but in the failure scenario Figure 5 (f), the arm remains extended after opening the door. Finally, Figure 5 (g) and (h) show the Close Sink skill, where the height of the blue water column visually correlates with the faucet handle angle. In Figure 5 (g), the faucet is successfully closed, while in Figure 5 (h), the arm accidentally collides with the handle, causing the opening angle to increase rather than decrease.
Figure 5: Additional Qualitative Results of Different Skills. (a), (b) Success and failure cases of the Pick skill. (c), (d) Success and failure cases of the Place skill. (e), (f) Success and failure cases of the Open Door skill. (g), (h) Success and failure cases of the Close Sink skill.
Appendix B Prompt for Task Planning
To ensure a fair comparison, we utilize a standardized prompt structure across all task planners. The input consists of two components: a system prompt and a user prompt. The system prompt defines the detailed rules and logic for task planning, while the user prompt provides the environmental context, including visual and acoustic observations.
Prompt for Task Planning in SonicStow and SonicInteract. The system prompt for SonicStow and SonicInteract is structured into four components. The first component defines the strictly allowed skill vocabulary. The second establishes the logical mapping between audio and the corresponding skill chains. The third specifies the planning constraints and the mandatory JSON output format. Finally, the fourth component provides a structural template for the response. Complementing this, the user prompt supplies the visual and acoustic observations to ground the planning process.
(I) Prompt for Task Planning in SonicStow and SonicInteract System Prompt: You are an expert Embodied AI Task Planner. Your goal is to generate a strictly ordered list of skills based on the input auditory (audio) and visual (head_rgb) observations. 1. ALLOWED SKILL VOCABULARY: You are strictly limited to ONLY the following skills. Do not invent new words or change spellings: [“nav”, “pick”, “place”, “open_door”, “close_sink”]. 2. LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound and classify the source object to determine the skill sequence: 1. Category A: SonicStow Objects (a) If the sound source is: Mechanical Alarm, Phone (ringing tone), or Furby (toy). REQUIRED PLAN: [“nav”, “pick”, “place”] 2. Category B: SonicInteract Objects (a) If the sound source is: Doorbell. REQUIRED PLAN: [“nav”, “open_door”] (b) If the sound source is: Water Running. REQUIRED PLAN: [“nav”, “close_sink”] 3. CONSTRAINTS and OUTPUT FORMAT: 1. Never return an empty list. 2. Output MUST be a valid JSON object containing a single key “plan”. 3. Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string. 4. OUTPUT FORMAT: {“plan”: [“skill_1”, “skill_2”, …]}
User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”
Prompt for Task Planning in Bi-Sonic Manipulation. For the Bi-Sonic Manipulation task, we organize the system prompt into five segments to handle the dual-source complexity. The initial segment lists the permissible skill vocabulary. The second details the correspondence between sound categories and their required skill sequences. Crucially, the third segment enforces interaction priority. We explicitly identify the primary sound source, requiring the planner to deduce the secondary source from the mixed audio and generate a sequential execution plan for both. The fourth and fifth segments impose strict output constraints and provide a valid JSON template to ensure parsing reliability. Finally, the user prompt provides the model with multimodal context, including visual frames and audio data.
(II) Prompt for Task Planning in Bi-Sonic Manipulation System Prompt: You are an expert Embodied AI Task Planner. Your goal is to generate a strictly ordered list of skills based on the input auditory observations. There are exactly two sound sources detected. 1. ALLOWED SKILL VOCABULARY (STRICT): You are strictly limited to ONLY the following skills: [“nav”, “pick”, “place”, “open_door”, “close_sink”]. Do not invent new words or change spellings. 2. LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound to determine the action sequence: 1. Category A: SonicStow Objects (a) Sources: “Mechanical_Alarm”, “Phone”, “Furby” $\rightarrow$ [“nav”, “pick”, “place”] 2. Category B: SonicInteract Objects (a) Source: “Doorbell” $\rightarrow$ REQUIRED PLAN: [“nav”, “open_door”] (b) Source: “Running-Water” $\rightarrow$ REQUIRED PLAN: [“nav”, “close_sink”] 3. OBJECT MAPPING INSTRUCTION (CRITICAL): You are provided with a partial hint about the audio content: One of the sound sources is CONFIRMED to be: ‘‘{obj_1}". You have to interact with it first. Your task is to identify the other sound source from the audio and generate the plan. You MUST map the objects exactly as follows: 1. Key “first_sound”: Generate the plan for ‘‘{obj_1}". 2. Key “second_sound”: Identify the SECOND sound source from the audio (the one that is NOT ‘‘{obj_1}") and generate its plan. 3. Do NOT output plans for any other potential objects unless they are detected as the second sound. 4. CONSTRAINTS and OUTPUT FORMAT: 1. Output MUST be a valid JSON object containing a single root key “plan”. 2. The keys inside “plan” MUST be exactly “first_sound” and “second_sound”. 3. Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string. 5. Expected JSON Structure: { “plan”: { “first_sound”: [“skill_1”, “skill_2”, …], # $\leftarrow$ Actions for the object assigned to ’first_sound’ “second_sound”: [“skill_1”, “skill_2”, …] # $\leftarrow$ Actions for the object assigned to ’second_sound’ } }
User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”

### v2

@@
 From Instruction to Event: Sound-Triggered Mobile Manipulation
-Abstract
+Abstract.
 Current mobile manipulation research predominantly follows an instruction-driven paradigm, where agents rely on predefined textual commands to execute tasks.
@@
 Notably, in the challenging dual-source scenario, the agent successfully isolates the primary source from overlapping acoustic interference to execute the first interaction, and subsequently proceeds to manipulate the secondary object, verifying the robustness of the baseline.
-Machine Learning, ICML
-Figure 1: Motivation for our work.
-(Upper) An example of sound-triggered mobile manipulation, the trigger signal is the sound of the event rather than instructions from humans.
-(Left) Instruction-based mobile manipulation needs humans to analyze the trigger signal and manually give the instruction for the robot model, which is passive.
-(Right) In sound-triggered mobile manipulation, the robot model automatically receives sound signals and actively outputs executable actions.
-1 Introduction
-Recent advances in Embodied AI have witnessed remarkable progress in mobile manipulation, enabling robots to navigate unstructured environments and interact with physical objects (Song et al., 2025a; Bai et al., 2025; Yu et al., 2025).
-Such synergy between mobility and interaction empowers agents to perform long-horizon household tasks such as object rearrangement (Ji et al., 2025; Huang et al., 2025; Escoriza et al., 2025) and multi-stage navigation (Song et al., 2025b; Lin et al., 2025), bringing us closer to general-purpose home robotic assistants.
-Current paradigm predominantly studies mobile manipulation under an instruction-conditioned setting, where the task goal and intended interaction are explicitly specified by a natural language instruction (Wu et al., 2025; Shah et al., 2025; Yan et al., 2025; Zhang et al., 2025).
-While effective for user-specified chores, this paradigm is ill-suited for the stochastic nature of daily life.
-For example, events such as a boiling kettle or a ringing doorbell typically happen unexpectedly and require timely intervention, where the latency inherent in human-in-the-loop interaction renders timely intervention impossible.
-This uncertainty is further compounded when multiple events occur concurrently or in rapid succession, i.e., waiting for a phone call from friends while expecting a doorstep delivery from a courier simultaneously.
-The indeterminate sequence of such events prevents the user from formulating a valid instruction plan in advance.
-This limitation underscores the discrepancy between static, predefined instructions and dynamic, sound-triggered scenarios (See Figure 1).
-Although recent efforts have attempted to generalize instructions ranging from concrete commands (Yenamandra et al., 2023) to abstract goals (Wu et al., 2023) and ambiguous queries (Ichter et al., 2022; Li et al., 2025), addressing these stochastic events remains an open challenge.
-In this context, we find that sound signals play a pivotal role.
-For instance, if the phone rings prior to a knock at the door, the robot should prioritize picking up the phone before retrieving the delivery.
-In such cases, robotic actions are modulated by stochastic sounds rather than predefined text instructions.
-Specifically, sounds serve a threefold function: (1) temporal trigger indicating the start of a stochastic event; (2) semantic cue guiding the robot’s action selection; and (3) spatial cue revealing the event’s location.
-Based on these insights, we introduce sound-triggered mobile manipulation, comprising SonicStow, SonicInteract, and Bi-Sonic Manipulation, where acoustic signals replace conventional textual instructions as the trigger signal.
-SonicStow and SonicInteract present unique challenges beyond traditional visual manipulation: (1) Acoustic semantic understanding, which requires the robot to interpret sounds to determine target locations and actions.
-(2) Sound-conditioned interaction, which demands the execution of articulated interactions solely triggered by audio.
-The Bi-Sonic Manipulation task further elevates the complexity by introducing dual sound sources, which demands the agent to prioritize the primary source amidst acoustic interference and subsequently locate and interact with the secondary sound source.
-To validate our proposed tasks, we introduce Habitat-Echo, a simulation platform that seamlessly integrates acoustic rendering with interactive physical environments.
-Built upon the visual and physical capabilities of Habitat (Szot et al., 2021; Puig et al., 2024), Habitat-Echo addresses the absence of sound signals in prior simulators.
-Specifically, we leverage room impulse responses to convolve source sound with the environment’s geometry, thereby ensuring physically accurate sound propagation and spatialization from the source to the receiver.
-Leveraging this platform, we construct a benchmark tailored for sound-triggered mobile manipulation.
-The dataset encompasses many interactive assets, categorized into: (1) static silent rigid objects, (2) sound-emitting rigid objects, and (3) sound-emitting articulated objects.
-We utilize this dataset to train policy models and further introduce an Omni-modal Large Language Model (Omni-LLM) as a high-level task planner to understand acoustic events and decompose complex manipulation goals.
-The main contributions of this work are summarized as follows:
-  • Novel Sound-Triggered Tasks.
-Distinct from traditional instruction-following paradigms, we formulate three novel manipulation tasks triggered by stochastic acoustic signals.
-In these tasks, robotic actions are autonomously triggered by stochastic sound events rather than static predefined instructions, better reflecting the unpredictable and dynamic nature of real-world household scenarios.
-  • New Audio-Visual-Physical Simulation Platform.
-We introduce Habitat-Echo, a platform that integrates acoustic and visual rendering with interactive physical environments.
-By incorporating sound-emitting objects, Habitat-Echo facilitates large-scale episode generation and closed-loop online interaction, establishing a robust foundation for future research in embodied audio-visual learning.
-  • Hierarchical Baseline and Validation.
-We develop a hierarchical framework that is composed of a high-level task planner to interpret acoustic cues and several low-level executors for precise control.
-We leverage an Omni-LLM as the task planner and policy models as executors.
-The baseline framework achieves promising performance on the proposed benchmark.
-Figure 2: Illustration of the proposed tasks.
-(a) SonicStow requires the agent to interact with the sound source (a rigid object) via Navigate, Pick, and Place skills.
-(b) SonicInteract requires the agent to interact with the sound source (an articulated object) via Navigate, Open Door, and Close Sink skills.
-(c) Bi-Sonic Manipulation requires the agent to interact with the sound source (two objects sequentially) via Navigate, Pick, Place, Open Door, and Close Sink skills.
-2 Related Work
-Sound Navigation.
-Sound Navigation, proposed by (Chen et al., 2020), requires an agent to navigate to an audio source without explicit goal coordinates.
-The primary challenges involve understanding physical spaces, localizing sound emitters, and avoiding obstacles.
-Existing methods can be broadly categorized into two streams: real-world application and versatile training.
-For real-world application, researchers aim to bridge the simulation-to-reality gap by introducing complex environmental settings.
-(Chen et al., 2021) assign semantic meanings to audio sources and introduce dynamic sound emission, replacing the continuous, generic signals used in prior work.
-(Younes et al., 2023) further extends the setting from static to moving sound sources.
-To simulate complex acoustic environments, (Yu et al., 2022) introduces an adversarial sound attacker that generates distracting noises with varying categories and volumes.
-For versatile training, methods leverage cross-modal synergies to enhance performance.
-VXN (Wang et al., 2022) unifies diverse navigation tasks, using sound navigation to facilitate the training of visual-language navigation.
-Recently, (Kondoh and Kanezaki, 2025) incorporates language descriptions of past and future actions as an auxiliary task to enhance policy learning.
-However, these works mainly focus on navigation capabilities, neglecting physical interactions with objects within the environment.
+Audio, Mobile Manipulation, Task Planning
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 10–14, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06submissionid: 7623ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Robotic planning
+Figure 1.
+Motivation for our work.
+This scenario presents two simultaneous and stochastic events triggered by sounds: a ringing doorbell and an alarm.
+The agent is required to actively explore the environment.
+Unlike the conventional instruction-driven paradigm, Sound-Triggered Mobile Manipulation (STMM) can effectively handle unpredictable events without relying on explicit textual instructions.
+1. Introduction
+Recent advances in embodied AI have led to remarkable progress in mobile manipulation, enabling robots to navigate in unstructured environments and physically interact with surrounding objects (Song et al., 2025a; Bai et al., 2025; Yu et al., 2025).
+Such synergy between mobility and interaction empowers agents to perform long-horizon household tasks such as object rearrangement, multi-stage interaction, and sequential task execution (Ji et al., 2025; Huang et al., 2025; Li et al., 2025b; Escoriza et al., 2025; Song et al., 2025b; Lin et al., 2025), pushing embodied agents closer to general-purpose household assistance.
+Despite these advances, most existing mobile manipulation works are built upon an instruction-driven paradigm, where the task goal is explicitly specified by the user through instructions (Wu et al., 2025; Shah et al., 2025; Yan et al., 2025; Zhang et al., 2025; Liu et al., 2025a; Li et al., 2025a), and the agent is expected to execute the instructed behavior accordingly.
+While this paradigm excels at completing user-specified chores, it also implicitly assumes that the task objective is known before execution begins.
+However, in real-world household environments, many tasks are triggered by stochastic events that require a timely response.
+For example, a ringing doorbell may prompt the agent to approach the door and open it, and a ringing alarm may indicate that the device needs to be turned off (see Figure 1).
+The unpredictable nature of these events makes it hard for users to monitor the situation and provide explicit instructions in time.
+This highlights the urgent need for event-driven mobile manipulation, where an intelligent agent should be able to proactively perceive and respond to stochastic environmental events.
+Among various environmental signals, sound serves as a natural trigger for event-driven interaction.
+Sound is omnidirectional and independent of illumination compared to visual cues.
+In fact, many of these events are exposed acoustically long before they enter the visual field.
+An agent can therefore detect critical events even when the sounding target lies outside its field of view or under poor lighting conditions.
+Moreover, sound signals not only indicate the occurrence of an event, but also provide partial semantic cues about the event type and spatial cues about the location of the sound source.
+These properties make sound a uniquely informative modality for proactive embodied response.
+Motivated by this observation, we introduce a novel problem termed Sound-Triggered Mobile Manipulation (STMM), where the agent must autonomously ground the sound source and execute appropriate interactions based on the auditory event and scene context.
+Unlike conventional instruction-driven mobile manipulation, STMM requires the agent to infer the task goal from environmental events rather than receiving it directly from a human user (see Figure 1).
+Concretely, we categorize sound-triggered mobile manipulation into two fundamental interaction types: (1) object relocation, where the agent localizes a sounding rigid object (e.g., a phone) and alters its spatial position via pick-and-place skills, and (2) state transition, where the agent localizes a sounding functional or articulated object (e.g., a door) and manipulates it to change its internal physical states, such as opening or closing.
+Building upon this, we investigate both single-event and more challenging multi-event settings.
+In the multi-event setting where multiple sound events occur simultaneously, the agent must not only determine the correct interaction for each event, but also resolve event prioritization and execute tasks in an appropriate order.
+Overall, sound-triggered mobile manipulation demands the agent to jointly reason about what event is happening, where the relevant sounding target is located, and how the correct interaction should be carried out, while additionally handling task scheduling in multi-event scenarios, making the problem particularly challenging.
+However, existing embodied simulators either focus strictly on visual-physical interaction (e.g., mobile manipulation simulator (Puig et al., 2024)) or visual-sound sensing (e.g., sound navigation environment (Chen et al., 2020)).
+There is a lack of an integrated environment where agents can perceive spatial sound and physically interact with sound-emitting targets within a closed loop, thus significantly hinders systematic research on sound-triggered mobile manipulation.
+To bridge this gap, we develop Habitat-Echo, a unified audio-visual-physical simulation platform.
+By seamlessly integrating spatial audio rendering into existing interactive embodied environments and extending it with abundant sound-emitting interactive assets, Habitat-Echo enables a closed-loop simulation of sound perception, sound source grounding, and physical interaction within a single unified platform.
+Built on Habitat-Echo, we further establish STMM-1.9K, the first benchmark for sound-triggered mobile manipulation, covering diverse sounding objects and interaction scenarios under both single-event and multi-event settings.
+To provide a first-step solution to this new problem, we design a hierarchical baseline framework that decomposes sound-triggered mobile manipulation into high-level acoustic reasoning and low-level physical execution.
+For high-level acoustic reasoning, a task planner powered by an Omni-MLLM (Xu et al., 2025) is leveraged to translate audio-visual observations into an executable skill chain, including task priority and the sequence of required skills.
+For low-level physical execution, a set of lightweight policy models is invoked sequentially to carry out each required skill.
+Each policy model focuses on a specific skill, such as navigation, pick-and-place, and door-opening, enabling specialized control for diverse interaction types.
+Experimental results on the STMM-1.9K benchmark provide an initial validation of the proposed setting and reveal its core challenges, including auditory grounding, task prioritization, and multi-stage execution.
+Our contributions are summarized as follows:
+  • We propose Sound-Triggered Mobile Manipulation (STMM), a novel event-driven embodied task in which agents are required to respond to sound events physically without relying on explicit per-episode instructions.
+  • We develop Habitat-Echo, a unified embodied simulation platform that seamlessly unifies high-fidelity spatial audio perception with interactive physical interaction.
+  • We establish STMM-1.9K, the first benchmark for STMM, and propose a hierarchical baseline framework that offers an initial solution and reveals the core challenges of this new embodied task.
+Figure 2.
+Different settings of Sound-Triggered Mobile Manipulation (STMM).
+(a) Object Relocation in the single-event setting requires the agent to interact with the sound source (a rigid object) via Navigate, Pick, and Place skills.
+(b) State Transition in the single-event setting requires the agent to interact with the sound source (an articulated object) via Navigate, Open Door, and Close Sink skills.
+(c) Multi-event Setting requires the agent to interact with the sound sources sequentially.
+2. Related Work
 Mobile Manipulation.
@@
 Relying on predefined textual commands limits the agent’s ability to automatically process and actively respond to dynamic events in the environment.
-Task Planning for Robotic Manipulation.
-Task planning bridges high-level reasoning with low-level execution.
+Sound in Embodied AI.
+Sound provides crucial spatial and contextual information for embodied agents.
+Existing applications of sound in embodied AI can be broadly categorized into two main areas: navigation and manipulation.
+Audio-Visual Navigation, pioneered by (Chen et al., 2020), requires an agent to navigate to an audio source without explicit goal coordinates.
+Existing methods generally follow two streams: complex environmental modeling for real-world applications and cross-modal strategies for versatile training.
+For sim-to-real gap, (Chen et al., 2021) replaces continuous, generic signals with dynamic, semantically meaningful sound emissions (Chen et al., 2021).
+This has been further extended to moving sound sources (Younes et al., 2023) and complex acoustic environments featuring adversarial attackers that generate distracting noises (Yu et al., 2022).
+For versatile training, VXN (Wang et al., 2022) unifies diverse navigation tasks by using sound navigation to facilitate vision-language training.
+(Kondoh and Kanezaki, 2025) incorporates language descriptions of past and future actions as an auxiliary task to enhance policy learning.
+However, these works predominantly focus on navigation capabilities, neglecting fine-grained physical interactions with objects.
+In robotic manipulation, sound typically serves as an informative modality to guide physical processes.
+(Wang et al., 2025a) utilizes MMAudio (Ruan et al., 2023) to generate the sound of pouring liquid into a bottle to assist the process.
+RoboOmni (Wang et al., 2026) extends textual instructions into audio contexts, enabling interactive communication with humans.
+Additionally, (Guo et al., 2026) employs beamforming to convert acoustic signals into 2D masks, complementing visual observations.
+Nevertheless, these approaches primarily treat sound as a complementary input, rather than the primary triggering signal that initiates an event.
+Task Planning.
+Recent LLM advancements greatly broaden the scope of task planning.
+Specifically, Tool-Planner (Liu et al., 2025b) minimizes error propagation in tool usage by effectively grouping and selecting APIs.
+In algorithmic planning, (Katz et al., 2024) boosts planning efficiency while maintaining logical rigor through optimized tree-based methods.
+For web agents, (Wang et al., 2025b) proposes a benchmark evaluating LLMs’ ability to translate user intentions into step-by-step product associations.
+Task planning in robotic manipulation bridges high-level reasoning with low-level execution.
 Early approaches (Kaelbling and Lozano-Pérez, 2011; Toussaint, 2015; Srivastava et al., 2014) primarily relied on predefined symbolic representations for continuous control.
 Recent research leverages the extensive world knowledge within LLMs and VLMs for planning, broadly categorized into two streams The first stream focuses on task decomposition.
-Wake et al.
-(Wake et al., 2024) utilize GPT-4V (OpenAI, 2023) to ground spatial-temporal information within the demonstration video and analyze affordances for plan generation.
-Similarly, Kachaev et al.
-(Kachaev et al., 2025) employ VLMs to decompose instructions into subtask sequences.
+Wake et al(Wake et al., 2024) utilize GPT-4V (OpenAI, 2023) to ground spatial-temporal information within the demonstration video and analyze affordances for plan generation.
 RoboBrain (Ji et al., 2025) unifies planning with affordance perception and trajectory prediction within a single model, enhancing manipulation capabilities.
 The second stream emphasizes feedback-based refinement.
-Zhi et al.
-(Zhi et al., 2025) optimize code-based plans using execution feedback, and Feng et al.
-(Feng et al., 2025) improve plans by anticipating future states.
-Similarly, Wang et al.
-(Wang et al., 2025) employ VLMs to solve PDDL problems constructed from current observations.
+Zhi et al(Zhi et al., 2025) optimize code-based plans using execution feedback, and Feng et al(Feng et al., 2025) improve plans by anticipating future states.
 However, these works primarily rely on visual and textual inputs, neglecting the importance of audio signals.
-3 Task Definition and Skill
-3.1 Task Definition
-We formally define three novel sound-triggered tasks: SonicStow, SonicInteract, and Bi-Sonic Manipulation.
-SonicStow and SonicInteract are mobile manipulation tasks with a single sound source.
-Bi-Sonic Manipulation, a scenario with dual sound sources.
-See Figure 2 for more details.
-SonicStow.
-This task simulates a scenario where an agent actively perceives an acoustic signal, localizes its source, and performs Pick and Place operations on household objects.
-Distinct from standard instruction-following paradigms that rely on predefined instructions, the agent must autonomously identify the target and conduct interaction solely based on the stochastic audio event.
-In each episode, the agent is initialized with a random pose and start location, while a single sound source, randomly selected from three objects (Phone, Alarm, Furby) within the scene (see Figure 2(a)).
-To introduce visual clutter, two distractor objects are randomly sampled from the YCB dataset for each episode.
-All objects are placed on elevated receptacles (e.g., existing furniture or inserted tables) instead of floor.
-SonicInteract.
-This task entails localizing an acoustic event and performing manipulation on articulated objects.
-Different from picking and placing objects in SonicStow, this task requires the agent to precisely control the end-effector and manipulate articulated mechanisms, i.e., opening a door or turning off a faucet in the sink (see Figure 2(b)).
-In each episode, the agent is initialized with a random pose and start location.
-The target single-sound source is randomly sampled from two categories: a doorbell or a sink with running water.
-Object placement follows category-specific constraints: doorbells are fixed to the existing doors within the scene, whereas sinks are placed on the floor with arbitrary orientations and locations.
-Bi-Sonic Manipulation.
-The Bi-Sonic Manipulation task integrates the challenges of SonicStow and SonicInteract, requiring the agent to manipulate both rigid and articulated objects (see Figure 2(c)).
-Additionally, this task introduces three more challenges.
-The first challenge of this task is the presence of two simultaneous sound sources, which creates a complex acoustic field featuring signal dynamics and superposition.
-Unlike single-source tasks, the agent must distinguish and interact with both sources sequentially based on a given priority (i.e., the main source followed by the secondary source).
-The categories and spatial locations of these sources are randomized.
-The second challenge of this task is two-time interactions.
-After interacting with the first object, the sound of the first object stops, and there is only one sound source in the scene.
-The agent has to execute skills and interact with the second object.
-Furthermore, to enhance visual complexity, we still introduce unrelated objects (Calli et al., 2015) to serve as visual clutter.
-3.2 Skill
-Similar to previous work (Gu et al., 2023; Shukla et al., 2024; Wang et al., 2025), the proposed tasks are decomposed into different subtasks.
-Each subtask is solved by an individual skill.
-A task is considered completed successfully if and only if all constituent skills are successfully executed in sequence.
-There are five skills in the proposed three tasks.
-Navigate.
-The action space consists of four actions: MoveForward, TurnLeft, TurnRight, and Stop.
-Specifically, MoveForward translates the agent by 0.5m along its local forward axis, while rotational actions (TurnLeft, TurnRight) execute $90^{\circ}$ in-place rotations to the left and right, respectively.
-The Stop action terminates the episode without changing the agent’s pose and location.
-An episode is considered successful if and only if the agent executes Stop while positioned within a region of 0.5m radius of the goal.
-Pick.
-The agent operates in an 8-DoF continuous action space.
-The first seven dimensions control the arm’s joint angles, driving the end-effector to the target pose via the kinematic chain (Abdelmaksoud et al., 2024), while the final dimension modulates the gripper state.
-A Pick skill is considered successful with two conditions: (1) Grasping: The gripper action is positive while an object is within a 15cm snapping threshold of the end-effector, triggering a stable grasp.
-(2) Retraction: The agent effectively holds the object and returns the arm to a predefined resting pose, ensuring a smooth transition to subsequent subtasks.
-Place.
-The agent operates in an 8-DoF action space identical to that of the Pick skill.
-The success conditions for the Place skill are twofold: (1) Releasement: The gripper opens to release the object, ensuring it settles within 0.15cm of the target position.
-(2) Retraction: Upon release, the arm reverts to the default resting pose, preparing for the next subtask.
-Open Door.
-Operating within the 8-DoF action space, the agent’s end-effector attaches to the handle of the door and drives it along its own functional trajectory.
-The skill is considered successful with two conditions: (1) Opening: While connected with the door handle, the agent must drive the door’s joint angle to a specified target threshold.
-(2) Retraction: The agent disengages from the door handle and returns the arm to the resting pose.
-Close Sink.
-Similar to the Open Door skill, the agent utilizes the 8-DoF action space to manipulate the sink’s rotary mechanism.
-The end-effector must grasp the faucet handle and apply torque to rotate it along its functional axis.
-Success is determined by two conditions: (1) Closure: While maintaining grasp, the agent must rotate the handle to the zero-degree threshold within a tolerance of 0.2 rad.
-(2) Retraction: Upon completing the rotation, the agent releases the handle and reverts the arm to the resting pose.
-Table 1: (a) Comparison of different simulators.
-Our Habitat-Echo supports sound rendering and physical interaction at the same time.
-(b) Distribution of audio assets across different categories in Habitat-Echo.
-We ensure a diverse split of sound instances for training and test sets.
+3. Sound-Triggered Mobile Manipulation
+3.1. Problem Formulation
+Unlike conventional instruction-driven mobile manipulation, which is typically formulated as learning a policy $\pi(a_{t+1}\mid O_{t},L)$ conditioned on an explicit instruction $L$, Sound-Triggered Mobile Manipulation (STMM) removes the explicit instruction that specifies the task goal.
+Instead, we formulate STMM as a partially observable Markov decision process.
+At each time step $t$, the agent only receives a multimodal observation $O_{t}=\{V_{t},A_{t},P_{t}\}$, where $V_{t}$, $A_{t}$, and $P_{t}$ denote visual observations, sound observations, and proprioceptive states, respectively.
+Triggered by environmental sound events $E_{sound}$, the agent must implicitly infer the task goal $\mathcal{G}$ from its observations and execute actions $a_{t}\in\mathcal{A}$ over time to interact with the relevant target.
+The episode is considered successful if the event-conditioned goal induced by $E_{\text{sound}}$ is successfully accomplished.
+Concretely, as shown in Figure 2, depending on the physical properties of the sound-emitting targets associated with $E_{sound}$, we categorize the required interactions to fulfill $\mathcal{G}$ into two fundamental types:
+Object Relocation.
+Here, the sounding target is a rigid object (e.g., a ringing phone or alarm in Figure 2(a)), and the goal is to resolve the event by changing its spatial position (e.g., retrieving the ringing phone).
+To accomplish this, agent must localize and approach the relevant sound source, and then perform object-centered manipulation such as grasping and placing.
+The main challenge lies in translating cross-modal grounding to executable object interaction.
+State Transition.
+Here, the sounding target is a functional or articulated object governed by object-specific kinematic constraints (e.g., a running faucet or a door associated with a ringing doorbell in Figure 2(b)), and the goal is to resolve the event by changing its internal physical state (e.g., turning of the faucet).
+Compared to object relocation, since the agent must operate along the object’s functional axis or articulated trajectory, this interaction type imposes stronger constraints on local geometry and manipulation precision.
+For example, the agent must precisely grasp a faucet handle and rotate it along its functional axis to stop the running water.
+3.2. Task Setting
+Beyond the interaction types defined above, we further design single-event and multi-event settings for STMM, which differ in the complexity of source localization and decision making.
+Single-event Setting.
+In the single-event setting, only one sounding event is present in the scene.
+The agent is required to infer the event-conditioned goal from the multimodal observation, localize the relevant target, and execute the corresponding interaction to resolve the event.
+Depending on the physical properties of the sounding target, the required interaction can take the form of either object relocation or state transition.
+Multi-event Setting.
+In the multi-event setting, multiple sound events occur simultaneously in the same scene, leading to a more complex acoustic field with signal superposition and source interference.
+Compared with the single-event setting, the agent must not only isolate the primary sound source from interference but also perform priority reasoning to execute the required interactions in an appropriate order.
+After the first event is resolved, its corresponding sound disappears and the acoustic scene changes, requiring the agent to continue interacting with the remaining target under an updated environment state.
+3.3. Skill Space
+In STMM, directly learning a policy from raw multimodal observations to low-level actions inextricably entangles high-level event reasoning with low-level physical execution, which imposes severe learning burdens and hinders proactive decision-making.
+To mitigate this, we decouple these complex embodied responses into a set of reusable skill primitives, where each skill represents an atomic sensorimotor mapping tailored to a specific physical interaction.
+Specifically, the skill space is define as:
+(1) $\mathcal{K}=\{\texttt{Navigate},\texttt{Pick},\texttt{Place},\texttt{OpenDoor},\texttt{CloseSink}\}.$
+Here, Navigate denotes approaching the relevant sounding target.
+Pick and Place correspond to rigid-object relocation via the agent’s gripper.
+OpenDoor and CloseSink correspond to state-transition interactions.
+The detailed execution spaces and termination criteria are elaborated in the Supplementary materials.
+Under this abstraction, different events can be solved through different skill compositions, i.e., the Skill Chain.
+An event is considered resolved successfully if and only if all constituent skills are successfully executed in sequence.
+Object relocation is typically instantiated as:
+(2) $\texttt{Navigate}\rightarrow\texttt{Pick}\rightarrow\texttt{Place}],$
+whereas state transition is instantiated as:
+(3) $\texttt{Navigate}\rightarrow\texttt{OpenDoor}]~~or~~[\texttt{Navigate}\rightarrow\texttt{CloseSink}].$
+For multi-event scenes, the complete task is further represented as the ordered composition of multiple skill chains with an explicit priority structure.
+In this way, the skill chain serves as a structured bridge from latent event interpretation to executable embodied responses.
+4. Platform and Benchmark
+4.1. Habitat-Echo Simulation Platform
+Table 1. Comparison of different simulators. Our Habitat-Echo supports sound rendering and physical interaction at the same time.
 Simulator Sound Rendering Physical Interaction
@@
 Habitat-Echo (Ours)
-Split Alarm Furby Phone Sink Doorbell Total
-Train 11 30 19 6 11 77
-Test 5 14 9 3 5 36
-Total 16 44 28 9 16 113
-4 Simulator and Benchmark
-4.1 Simulator: Habitat-Echo
-Existing simulators (Kolve et al., 2017; Makoviychuk et al., 2021; Szot et al., 2021) prioritize physical interaction with visual observations, while others (Chen et al., 2020, 2022) incorporate audio-visual rendering specifically for navigation tasks (see Table 1(a)).
-There remains an absence of simulators that integrate sound rendering with physical interaction.
-We address this limitation by introducing Habitat-Echo, an extension of Habitat (Szot et al., 2021) that enables sound rendering alongside physical interaction.
-To enable high-fidelity acoustic simulation within the environment, we leverage precomputed Room Impulse Responses (RIRs) in SoundSpaces (Chen et al., 2020).
+As shown in  Table 1, existing simulators either prioritize visual-physical manipulation  (Kolve et al., 2017; Makoviychuk et al., 2021; Szot et al., 2021) prioritize physical interaction with visual observations, or strictly focus on audio-visual navigation (Chen et al., 2020, 2022), failing to support the audio-visual-physical closed-loop interaction.
+To bridge this gap, we develop Habitat-Echo.
+Environment and Asset Composition.
+Habitat-Echo is built upon interactive indoor 3D environments (Szot et al., 2021) and extends them with a sound-emitting interactive asset library tailored for STMM.
+This library comprehensively covers the predefined interaction types, including graspable sounding rigid bodies (e.g., phones, alarms) and articulated or functional objects (e.g., doors, sinks with a faucet on), providing a robust asset foundation for both object relocation and state transition tasks.
+These assets are explicitly designed to act as both acoustic event sources and manipulation targets, thereby closing the loop between perception and action.
+Audio-Physical Simulation.
+To enable high-fidelity acoustic rendering, we leverage precomputed Room Impulse Responses (RIRs) in SoundSpaces (Chen et al., 2020).
 An RIR serves as an accurate transfer function characterizing the multipath propagation between a source and a receiver (Kuttruff, 2016).
@@
 With the help of RIRs, Habitat-Echo synthesizes the auditory observation by performing time-domain convolution between the source waveform and the corresponding RIR on the fly.
-Furthermore, exploiting the property of the acoustic field, our simulator supports simultaneous multi-source rendering by superimposing the convolved signals from distinct sources, which is essential for the Bi-Sonic Manipulation task.
+Furthermore, exploiting the property of the acoustic field, our simulator supports simultaneous multi-source rendering by superimposing the convolved signals from distinct sources, which is essential for the multi-event setting.
 The final output is rendered as binaural audio, providing the agent with realistic spatial cues analogous to human auditory perception.
-4.2 Benchmark
-Sounding Objects and Audio Assets.
-In addition to standard YCB objects (Calli et al., 2015), we introduce five more categories of interactive sounding objects.
-For SonicStow, we introduce graspable rigid objects, including phone, alarm, and furby.
-These objects can be manipulated via Pick and Place skills.
-For SonicInteract, we incorporate articulated objects, including the sink and door.
+Skill Interface.
+To support the skill space $\mathcal{K}$ defined in  Section.
+3.3, Habitat-Echo provides standardized execution interfaces and underlying observation/action space support for each skill primitive (e.g., supporting Navigate termination based on audio-visual features, or OpenDoor operations bounded by joint constraints).
+The platform allows for the independent initialization, termination, and success detection of each skill, thereby enabling the sequential scheduling and closed-loop evaluation of entire skill chains.
+4.2. STMM-1.9K Benchmark
+Figure 3.
+(a) Distribution of audio assets.
+The semantic are shown in different colors, ranging from Electronic Doorbell Alert to Ambient Notifications.
+(b) Episode statistics.
+Train and test splits.
+O.R.
+denotes Object Relocation, and S.T.
+denotes State Transition.
+Dataset Construction.
+In addition to YCB objects (Calli et al., 2015), we introduce more categories of interactive sounding objects.
+For object relocation, we introduce graspable rigid objects, including phone, alarm, and furby.
+For state transition, we incorporate articulated objects, including the sink and door.
 Unlike rigid items, these objects possess kinematic joints and are constrained to move along fixed functional axes.
 For the sink, we incorporate a dynamic water flow that is physically coupled to the faucet handle, moving synchronously with the handle’s rotation.
-To foster robust acoustic generalization, we also collect a diverse library of real-world sounds, implementing a strict one-to-many mapping where each object category is associated with multiple unique sound instances (see Table 1(b)).
-This design forces the agent to learn semantically meaningful audio-visual associations rather than overfitting to specific frequency patterns or waveforms, thereby ensuring high acoustic diversity of the benchmark.
+To foster robust acoustic generalization, we also collect a diverse library of real-world sounds, implementing a strict one-to-many mapping where each object category is associated with multiple unique semantics (see Figure 3(a)).
+This design forces the agent to learn semantically meaningful audio-visual associations rather than overfitting to specific frequency patterns or waveforms, ensuring high acoustic diversity of the benchmark.
 Episode Generation.
@@
 We also validate the solvability of every episode by ensuring a traversable path connects the start and goal, and we strictly verify that interactive objects have accessible frontal regions for valid manipulation, especially for sinks and doors.
-An episode is considered successful if and only if the required skills are successfully executed in sequence.
-5 Baseline
-Figure 3: Overview of the proposed baseline.
-The sound-triggered task planner processes initial audio-visual observations to reason and generate a high-level skill chain from the skill library.
-Guided by this chain, specialized policy models are sequentially activated to generate low-level actions and interact with the environment.
+In the single-event setting, the dataset provides 660 training and 222 testing episodes for both object relocation and state transition.
+The multi-event setting contains 660 training and 355 testing episodes (See Figure 3(b)).
+An episode is considered successful strictly when task planning succeeds, and all prerequisite skills are completed in the correct order.
+Consequently, performance in the proposed setting is measured by success rates of task planning and skill completion.
+Figure 4.
+Overview of the proposed baseline.
+The sound-triggered task planner processes the initial observation to reason and generate a high-level skill chain from the skill library (right).
+Guided by this chain, specialized policy models are sequentially activated to generate low-level actions and interact with Habitat-Echo.
+5. Hierarchical Baseline for STMM
 The proposed baseline comprises two core components, i.e., sound-triggered task Planner and policy models (see Figure 4).
@@
 Following this chain, the policy models serve as the low-level executors, sequentially translating each assigned skill into concrete actions to complete the corresponding subtasks.
-5.1 Sound-triggered Task Planner
+5.1. Sound-triggered Task Planner
 Motivated by MCP (Anthropic, 2024), we propose a sound-triggered task planner that operates as the high-level reasoning module.
@@
 This chain acts as a high-level directive, guiding the sequential activation of the following policy models to complete the task.
-5.2 Policy Model
+5.2. Skill-Specific Policy Model
 Policy models serve as low-level executors, translating the skill chain from the task planner into concrete actions and sequentially interacting with the environment.
@@
 The temporal module and actor-critic modules maintain a consistent structure following prior works (Chen et al., 2020; Gu et al., 2023), and the modality encoders are tailored to the specific observation space of different skills.
-For the Navigate skill, the observation space includes audio and head RGB, and the modality encoders are designed based on task complexity.
-In SonicStow and SonicInteract, the audio observations are converted into spectrograms and processed by a vanilla CNN encoder (Chen et al., 2020).
-In Bi-Sonic Manipulation, audio observations are transformed into Mel-spectrograms, and the encoder is LanguageBind (Zhu et al., 2024).
+For the Navigate skill, the observation space includes audio and head RGB, and the modality encoders (Chen et al., 2020; Oquab et al., 2024) are designed based on task complexity.
+In object relocation and state transition, the audio observations are converted into binaural spectrograms and processed by a vanilla CNN encoder (Chen et al., 2020).
+In the multi-event setting, audio observations are transformed into Mel-spectrograms, and the encoder is LanguageBind (Zhu et al., 2024).
 An additional text encoder is introduced to specify the target sound source.
 For manipulation skills, including Pick, Place, Open Door, and Close Sink, the observation spaces consist of head and arm depth, which are processed by a standard CNN encoder (Chen et al., 2020).
-Table 2: Results on the Habitat-Echo benchmark.
-We report the success rate ($\%$) of task planning, individual skills, and overall performance across different tasks.
-The higher the better.
-(a) Evaluation of SonicStow.
-(b) Evaluation of SonicInteract.
-(c) Evaluation of Bi-Sonic Manipulation.
-Success rates are reported in the format “First Sound Source / Second Sound Source”.
-Baseline Task Planning Navigate Pick Place Overall
-Individual N/A 57.66 88.73 70.27 N/A
-Oracle 100.0 57.66 86.72 81.98 40.99
-Qwen2.5-Omni-3B 12.16 55.56 93.33 71.43 4.50
-Qwen3-Omni-30B 68.92 55.56 87.06 72.97 24.33
-Qwen2.5-Omni-7B 78.38 51.72 85.56 79.22 27.48
-Method Task Planning Navigate Open Door Close Sink Overall
-Individual N/A 45.50 62.16 77.67 N/A
-Oracle 100.00 45.50 50.00 72.63 27.90
-Qwen2.5-Omni-3B 44.14 3.00 50.00 0.00 0.33
-Qwen3-Omni-30B 59.01 44.70 33.33 73.20 14.05
-Qwen2.5-Omni-7B 71.17 44.63 50.00 74.16 19.72
-Method Task Planning Navigate Pick Place Open Door Close Sink Overall
-Individual N/A 9.58 / 56.61 85.16 / 17.39 32.48 / 6.00 67.12 / 58.66 42.87 / 38.00 N/A
-Oracle 100.00 9.58 / 46.67 92.31 / 0.00 25.00 / 0.00 72.73 / 0.00 40.00 / 100.00 4.23 / 0.56
-Qwen2.5-Omni-3B 14.08 8.00 / 0.00 100.00 / 0.00 0.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.56 / 0.00
-Qwen2.5-Omni-7B 17.18 14.75 / 100.00 100.00 / 0.00 0.00 / 0.00 100.00 / 0.00 0.00 / 100.00 0.56 / 0.56
-Qwen3-Omni-30B 51.83 11.96 / 0.00 100.00 / 0.00 0.00 / 0.00 66.67 / 0.00 37.50 / 0.00 2.53 / 0.00
-Figure 4: Qualitative Visualization of Task Execution.
-We present the execution trajectories for (a) SonicStow, (b) SonicInteract, and (c) Bi-Sonic Manipulation.
-For each task, the left panel illustrates the top-down view of the environment, highlighting the agent’s trajectories and the location of sound sources.
-The right panel displays the corresponding sequence of third-person keyframes during the execution.
-Note that Bi-Sonic Manipulation involves navigating to and interacting with two distinct sound sources sequentially.
-6 Experiment
-6.1 Implementation
+6. Experiment
+6.1. Implementation
 Implementation Detail.
@@
 For each task, the dataset is split into 660 instances for training.
-The testing dataset comprises 222 instances for both SonicStow and SonicInteract, and 355 instances for Bi-Sonic Manipulation.
+The testing dataset comprises 222 instances for both Object Relocation and State Transition, and 355 instances for Multi-event Setting.
 Baseline.
-We train low-level policy models individually and then compose them with varying task planners, implementing three types of baselines.
-First, the Individual baseline evaluates each skill in isolation.
-This baseline bypasses the task planner to measure each skill without the influence of cascading errors.
-Second, the Oracle baseline employs a privileged task planner that provides the ground-truth skill chains, serving as a performance upper bound for the task planning.
-Finally, we integrate mainstream Omni-LLMs as task planners to evaluate audio-visual reasoning capabilities, including Qwen2.5-Omni (3B, 7B) (Xu et al., 2025) and Qwen3-Omni-30B-Instruct (Yang et al., 2025).
+We train low-level policy models individually and then compose them with varying task planners, implementing two types of baselines.
+First, the Oracle baseline employs a privileged task planner that provides the ground-truth skill chains, serving as a performance upper bound.
+Second, we integrate mainstream Omni-MLLMs as task planners to evaluate audio-visual reasoning capabilities, including Qwen2.5-Omni (3B, 7B) (Xu et al., 2025) and Qwen3-Omni-30B-Instruct (Yang et al., 2025).
 These task planners are all training-free.
 We refer to these baselines directly by their own names.
-6.2 Result
-Quantitative Result of SonicStow.
-We present the evaluation results on the SonicStow task in Table 2(a).
-The Oracle baseline, which utilizes ground-truth planning, establishes the performance upper bound with an overall success rate of 40.99%.
-The Individual baseline reflects the standalone success rate of each policy model (e.g., 88.73% for Pick), decoupling low-level execution and high-level planning.
-Regarding task planning, Qwen2.5-Omni-7B achieves the highest accuracy of 78.38%, surpassing both the smaller Qwen2.5-Omni-3B and the larger Qwen3-Omni-30B.
-Driven by this, Qwen2.5-Omni-7B generates the highest number of correct skill chains, thereby attaining the best overall execution performance of 27.48% among all task planners.
-Quantitative Result of SonicInteract.
-The evaluation results on the SonicInteract are shown in Table 2(b).
-With ground-truth task planning, the Oracle baseline achieves a 27.90% overall success rate, which is the theoretical upper bound of the policy models.
+Method Single-event setting Multi-event Setting
+State Transition Object Relocation
+Planning ($\uparrow$) Completion ($\uparrow$) Planning ($\uparrow$) Completion ($\uparrow$) Planning ($\uparrow$) Completion ($\uparrow$)
+Oracle 100.00 27.90 100.00 40.99 100.00 0.56
+Oracle* 100.00 81.08 100.00 70.72 100.00 37.18
+Qwen2.5-Omni-3B 44.14 0.33 12.16 4.50 14.08 0.00
+Qwen3-Omni-30B 59.01 14.05 68.92 24.33 51.83 0.00
+Qwen2.5-Omni-7B 71.17 19.72 78.38 27.48 17.18 0.56
+Qwen2.5-Omni-3B* 44.14 29.27 12.16 8.55 14.08 3.09
+Qwen3-Omni-30B* 59.01 53.15 68.92 48.19 51.83 19.71
+Qwen2.5-Omni-7B* 71.17 59.00 78.38 53.60 17.18 4.78
+Figure 5.
+Results on STMM-1.9k.
+(a) Success rates of task planning and skill completion.
+Qwen2.5-Omni-7B and Qwen3-Omni-30B receive the best performance on single-event and multi-event settings, respectively.
+(b, c) Progressive success rate of state transition completion.
+(d) Progressive success rate of object relocation completion.
+(e) Progressive success rate of skill completion in the multi-event setting.
+denotes the target position that is provided for the navigate policy model.
+Figure 6.
+Qualitative Visualizations of Task Execution.
+We present the execution trajectories for (a) Object Relocation, (b) State Transition, and (c) Multi-event Setting.
+For each task, left panel illustrates the top-down view, highlighting the agent’s trajectories and the location of sound sources.
+The right panel displays the sequence of third-person keyframes.
+Note that multi-event setting involves navigating to and interacting with two distinct sound sources sequentially.
+6.2. Result
+Quantitative Result of State Transition in Single-event setting.
+The evaluation results on the State Transition are shown in Figure 5 (a-c).
+Figure 5(a) shows the numerical performance of task planning and skill completion.
+Figure 5(b) and Figure 5(c) visualize the progressive success rate of State Transition.
+denotes the target position that is provided for the navigate policy model.
+With ground-truth task planning, the Oracle baseline achieves an 81.08% overall success rate, which is the theoretical upper bound of the policy models.
 Among the evaluated task planners, Qwen2.5-Omni-7B also emerges as the most effective one.
 It attains a planning rate of 71.17%, outperforming both Qwen2.5-Omni-3B and Qwen3-Omni-30B.
-The success rate of planning directly translates to execution success, and Qwen2.5-Omni-7B achieves the top overall performance of 19.72%.
+The success rate of planning directly translates to execution success, and Qwen2.5-Omni-7B and Qwen2.5-Omni-7B achieve the top overall performance of 59.00% and 19.72%, respectively.
 Conversely, we observe unsatisfactory cases in Qwen2.5-Omni-3B during the Close Sink subtask.
-Instead of responding to the sound of running water with the appropriate skill, Qwen2.5-Omni-3B consistently hallucinates a Pick and Place sequence, resulting in a near-zero success rate for this specific task.
-Quantitative Result of Bi-Sonic Manipulation.
-Table 2(c) presents the results for the Bi-Sonic Manipulation task.
-In this dual-source environment, Qwen3-Omni-30B achieves the highest task planning accuracy, showing superior robustness against acoustic distraction.
+Instead of responding to the sound of running water with the appropriate skill, Qwen2.5-Omni-3B consistently hallucinates a Pick and Place sequence, resulting in a near-zero success rate for this specific setting.
+Quantitative Result of Object Relocation in Single-event setting.
+We present the evaluation results on the Object Relocation in Figure 5 (a, d).
+The Oracle baseline, which utilizes ground-truth planning and target position, establishes the performance upper bound with an overall success rate of 70.72%.
+Regarding task planning, Qwen2.5-Omni-7B achieves the highest accuracy of 78.38%, surpassing both the smaller Qwen2.5-Omni-3B and the larger Qwen3-Omni-30B.
+Driven by this, Qwen2.5-Omni-7B generates the highest number of correct skill chains, thereby attaining the best skill completion performance of 53.60% among all.
+Quantitative Result of Multi-event Setting.
+Figure 5 (a, e) presents the results for the multi-event setting.
+In this multi-event environment, Qwen3-Omni-30B achieves the highest task planning accuracy, showing superior robustness against acoustic distraction.
 In contrast, Qwen2.5-Omni-7B fails to maintain its dominance in the task planning, receiving performance comparable to Qwen2.5-Omni-3B.
-Regarding the overall success rate, the Oracle and Qwen3-Omni-30B secure the top two results for the first sound source interaction.
-For the secondary source interaction, only the Oracle and Qwen2.5-Omni-7B manage to achieve 2 successful episodes.
-This near-zero success rate highlights the significant challenges posed by the Bi-Sonic Manipulation task (see Section.
-3.1).
-Table 3: Ablation Study on Input Modalities for Different Policy Models.
-We investigate optimal input combinations across three representative policy models for different skills, including Navigate, Pick, and Open Door.
-SR denotes success rate.
-(a) Policy model for Navigate.
-We evaluate different modality encoders including DINOv2 (Oquab et al., 2024) and CNN encoder (Chen et al., 2020) for visual inputs.
-Results indicate that combining Head RGB and Head Depth yields the highest success rate.
-(b) Policy model for Pick.
-Using a same encoder, the combination of Head Depth and Arm Depth outperforms single-view inputs.
-(c) Policy model for Open Door.
-Similarly, integrating Head Depth and Arm Depth is also proven effective for achieving optimal performance in the Open Door skill.
-Model Input Modality Modality Encoder SR ($\uparrow$)
-Head RGB Head Depth DINOv2 CNN
-I 57.66
-II 58.01
-III 51.80
-IV 61.71
-Model Input Modality SR ($\uparrow$)
-Head Depth Arm Depth
-I 57.20
-II 66.66
-III 88.73
-Model Input Modality SR ($\uparrow$)
-Head Depth Arm Depth
-I 46.39
-II 47.29
-III 62.16
+For skill completion, Oracle sets an upper bound success rate of 37.18%, with Qwen3-Omni-30B achieving 19.71%.
+Figure 5(e) shows the progressive success rate.
+Interference from the concurrent sound event poses a major challenge to completion of this setting (see Section.
+3.2).
 Qualitative Result.
@@
 In each subfigure, the left panel depicts the top-down trajectory and sound source locations, while the right panel displays sequential third-person keyframes of the interaction.
-Figure 4 (a) shows the SonicStow task.
+Figure 6 (a) shows the object relocation.
 Initially, the target object (an alarm) is located outside the agent’s field of view.
 Relying on sound signals for localization, the agent successfully navigates to the target and executes the Pick and Place skills.
-Figure 4 (b) illustrates the SonicInteract task.
+Figure 6 (b) illustrates the state transition.
 The agent accurately navigates to the sound-emitting door and performs the Open Door skill.
-Figure 4 (c) showcases the Bi-Sonic Manipulation task, which involves two simultaneous sound sources: a primary source (yellow) and a secondary source (green).
+Figure 6 (c) showcases the multi-event setting, which involves two simultaneous sound sources: a primary source (yellow) and a secondary source (green).
 Guided by the task planner, the agent prioritizes the primary source despite the acoustic interference.
@@
 More visualizations can be seen in supp.
-6.3 Ablation Study and Further Discussion
-Effect of Input Modality for the Navigate Skill.
-We evaluate the impact of different visual modalities, including head RGB, head depth, and their corresponding encoders for the Navigate skill (see Table 3(a)).
-First, the model relying solely on head depth with a CNN encoder yields the lowest performance.
+Table 2. Effect of audio modality on navigate policy model. Text Cat. and Tgt. Pos. denote the text category and target position of the sound source, respectively.
+Model Input Modality S.R. ($\uparrow$)
+Audio Text Cat. Tgt. Pos.
+I 0.00
+II 0.00
+III 57.66
+IV 94.59
+V 95.05
+Table 3.
+Ablation Study on Input Modalities.
+We investigate the impact of input modalities and identify optimal combinations across task planning and three representative policy models.
+S.R.
+denotes success rate.
+(a) Task Planning: Investigates the impact of Head RGB and sound, finding that combining two modalities is the best.
+(b) Policy model for Navigate (Visual Modality): Using Head RGB and Head Depth yields the highest result.
+(c) Policy model for Pick and Open Door: Combining Head Depth and Arm Depth outperforms single-view inputs for both Pick and Open Door.
+Method Input Modality S.R. ($\uparrow$)
+Head RGB Sound
+I 18.59
+II 51.55
+III 51.83
+Model Input Modality S.R. ($\uparrow$)
+Head RGB Head Depth
+I 51.80
+II 57.66
+III 61.71
+Model Input Modality S.R. ($\uparrow$)
+Head Depth Arm Depth Pick Open Door
+I 57.20 47.29
+II 66.66 46.39
+III 88.73 62.16
+6.3. Ablation Study and Further Discussion
+Effect of Audio Modality for Navigate Skill.
+Table 2 presents the impact of different input modalities for the navigate skill.
+Audio signals are converted into binaural spectrograms to provide implicit position information.
+Models lacking any spatial information (Model I and Model II) completely fail to navigate, yielding a 0.00% success rate.
+Relying solely on audio signals achieves a 57.66% success rate (Model III).
+Supplying the explicit target position directly reaches 94.59% (Model IV).
+Combining both audio and explicit target positions yields the best performance at 95.05% (Model V).
+Overall, these results indicate two key findings.
+First, spatial information of the sound source is crucial.
+Second, as an environmental signal, audio provides effective position information and is complementary to the explicit target position.
+Effect of input Modality for task planning.
+Table 3(a) shows the influence of various input modalities for task planning in the multi-event setting.
+We utilize Qwen3-Omni-30B (Yang et al., 2025) as the baseline.
+Relying exclusively on Head RGB yields the lowest success rate at 18.59%.
+Utilizing only sound signals substantially elevates the performance to 51.55%.
+Integrating both Head RGB and sound attains the optimal result of 51.83%.
+Consequently, these findings show the importance of auditory information in task planning, while visual input from Head RGB serves a supplementary function.
+Effect of Visual Modality for Navigate Skill.
+We evaluate the impact of different visual modalities, including head RGB, head depth for the Navigate skill (see Table 3(b)).
+First, the model relying solely on head depth yields the lowest performance.
 This underperformance arises because depth observations lack semantic texture, making it difficult for the model to learn meaningful associations between visual cues in depth observations and the semantic information contained in audio observations.
-Second, models using head RGB achieve better results, with the trainable CNN encoder performing slightly better than the frozen DINOv2 due to its training process that helps the model to adapt to the indoor environment.
+Second, models using Head RGB achieve better results.
 Third, integrating both Head RGB and Head depth yields the best overall performance.
 This combination proves effective because the modalities are complementary.
-Head RGB provides semantic context to align with audio sources, while depth offers precise geometric distance information.
-Despite the performance of the dual-modality method, we select head RGB with the DINOv2 encoder as our default configuration to maintain an optimal balance between performance and training efficiency.
-Effect of Input Modality for the Pick Skill.
-We investigate the impact of depth modalities on the Pick skill in Table 3(b), comparing the efficacy of head depth and arm depth.
-Using head depth alone yields the lowest success rate of 57.20%, primarily because the target object often falls outside the head camera’s optimal viewing range during manipulation.
+RGB provides semantic context to align with audio sources, while depth offers precise geometric distance information.
+Despite the performance of the dual-modality method, processing depth requires an additional visual encoder, which introduces additional computational burden.
+Therefore, we select single-modality Head RGB as the default configuration to maintain an optimal balance between performance, training efficiency, and computational cost.
+Effect of Input Modality for Pick Skill and Open Door Skill.
+We investigate the impact of depth modalities on the Pick skill and Open Door Skill in Table 3(c), comparing the efficacy of head depth and arm depth.
+For the Pick skill, using head depth alone yields the lowest success rate of 57.20%, primarily because the target object often falls outside the head camera’s optimal viewing range during manipulation.
 In contrast, the model relying solely on arm depth achieves a higher success rate of 66.66%, as the arm depth can directly observe the target when aligned.
 We observe that integrating both depth sources results in the best performance (88.73%).
-The two views prove complementary: the arm depth provides precise object-to-gripper distance, while the head depth offers global context to assist the arm in returning to the resting pose.
-Effect of Input Modality for the Open Door Skill.
-For the Open Door skill, we validate the impact of different depth inputs, specifically analyzing head depth and arm depth (see Table 3(c)).
-Models relying on a single depth source achieve comparable performance, with the head depth only model (Model II) slightly outperforming the arm depth only model (Model I).
-We attribute this margin to the fact that head depth provides a global view of the door’s structure, whereas the arm camera often suffers from initial misalignment with the door’s specific affordance (i.e., the handle).
-Crucially, combining both modalities (Model III) results in a dramatic performance boost to 62.16%.
-This improvement stems from the complementary nature of the inputs: the head depth offers the global context necessary to locate the door frame, which in turn allows the agent to leverage the arm depth for precisely perceiving the target affordance for accurate manipulation.
-7 Conclusion
+A similar trend appears in the Open Door Skill, where combining both depth views reaches the optimal result of 62.16%, significantly outperforming either single-modality setup.
+The two views prove complementary: the arm depth provides precise object-to-gripper distance and local affordances, while the head depth offers a global context of the gripper and object.
+7. Conclusion
 In this work, we introduce sound-triggered mobile manipulation, a novel paradigm that shifts agents from passive instruction-following to active, sound-driven interaction.
@@
 Extensive experiments validate that the proposed baseline enables agents to actively localize and manipulate sound-emitting objects.
-Notably, in challenging dual-source scenarios, the baseline demonstrates the capability to interact with the primary source amidst acoustic interference from a secondary signal, and successfully proceed to manipulate the subsequent target.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Additional Qualitative Result
-We present qualitative visualization results for the four manipulation skills in Figure 5.
-Figure 5 (a) and (b) illustrate the Pick skill.
-In the success case (a), the agent grasps the object and correctly returns the arm to the resting position, while in the failure case (b), the arm fails to retract to the standard pose despite holding the object.
-The Place skill is depicted in Figure 5 (c) and (d).
-In Figure 5 (c), the object is stably placed on the table, whereas in (d), the object is released with an unstable pose and falls off the surface of the small table.
-Figure 5 (e) and (f) visualize the Open Door skill.
-In the success scenario (e), the arm pulls the door open and resets to the resting position, but in the failure scenario Figure 5 (f), the arm remains extended after opening the door.
-Finally, Figure 5 (g) and (h) show the Close Sink skill, where the height of the blue water column visually correlates with the faucet handle angle.
-In Figure 5 (g), the faucet is successfully closed, while in Figure 5 (h), the arm accidentally collides with the handle, causing the opening angle to increase rather than decrease.
-Figure 5: Additional Qualitative Results of Different Skills.
-(a), (b) Success and failure cases of the Pick skill.
-(c), (d) Success and failure cases of the Place skill.
-(e), (f) Success and failure cases of the Open Door skill.
-(g), (h) Success and failure cases of the Close Sink skill.
-Appendix B Prompt for Task Planning
-To ensure a fair comparison, we utilize a standardized prompt structure across all task planners.
-The input consists of two components: a system prompt and a user prompt.
-The system prompt defines the detailed rules and logic for task planning, while the user prompt provides the environmental context, including visual and acoustic observations.
-Prompt for Task Planning in SonicStow and SonicInteract.
-The system prompt for SonicStow and SonicInteract is structured into four components.
-The first component defines the strictly allowed skill vocabulary.
-The second establishes the logical mapping between audio and the corresponding skill chains.
-The third specifies the planning constraints and the mandatory JSON output format.
-Finally, the fourth component provides a structural template for the response.
-Complementing this, the user prompt supplies the visual and acoustic observations to ground the planning process.
-(I) Prompt for Task Planning in SonicStow and SonicInteract System Prompt: You are an expert Embodied AI Task Planner.
-Your goal is to generate a strictly ordered list of skills based on the input auditory (audio) and visual (head_rgb) observations.
-1.
-ALLOWED SKILL VOCABULARY: You are strictly limited to ONLY the following skills.
-Do not invent new words or change spellings: [“nav”, “pick”, “place”, “open_door”, “close_sink”].
-2.
-LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound and classify the source object to determine the skill sequence: 1.
-Category A: SonicStow Objects (a) If the sound source is: Mechanical Alarm, Phone (ringing tone), or Furby (toy).
-REQUIRED PLAN: [“nav”, “pick”, “place”] 2.
-Category B: SonicInteract Objects (a) If the sound source is: Doorbell.
-REQUIRED PLAN: [“nav”, “open_door”] (b) If the sound source is: Water Running.
-REQUIRED PLAN: [“nav”, “close_sink”] 3.
-CONSTRAINTS and OUTPUT FORMAT: 1.
-Never return an empty list.
-2.
-Output MUST be a valid JSON object containing a single key “plan”.
-3.
-Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string.
-4.
-OUTPUT FORMAT: {“plan”: [“skill_1”, “skill_2”, …]}
-User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”
-Prompt for Task Planning in Bi-Sonic Manipulation.
-For the Bi-Sonic Manipulation task, we organize the system prompt into five segments to handle the dual-source complexity.
-The initial segment lists the permissible skill vocabulary.
-The second details the correspondence between sound categories and their required skill sequences.
-Crucially, the third segment enforces interaction priority.
-We explicitly identify the primary sound source, requiring the planner to deduce the secondary source from the mixed audio and generate a sequential execution plan for both.
-The fourth and fifth segments impose strict output constraints and provide a valid JSON template to ensure parsing reliability.
-Finally, the user prompt provides the model with multimodal context, including visual frames and audio data.
-(II) Prompt for Task Planning in Bi-Sonic Manipulation System Prompt: You are an expert Embodied AI Task Planner.
-Your goal is to generate a strictly ordered list of skills based on the input auditory observations.
-There are exactly two sound sources detected.
-1.
-ALLOWED SKILL VOCABULARY (STRICT): You are strictly limited to ONLY the following skills: [“nav”, “pick”, “place”, “open_door”, “close_sink”].
-Do not invent new words or change spellings.
-2.
-LOGIC MAPPING (AUDIO $\rightarrow$ SKILL): Analyze the input audio sound to determine the action sequence: 1.
-Category A: SonicStow Objects (a) Sources: “Mechanical_Alarm”, “Phone”, “Furby” $\rightarrow$ [“nav”, “pick”, “place”] 2.
-Category B: SonicInteract Objects (a) Source: “Doorbell” $\rightarrow$ REQUIRED PLAN: [“nav”, “open_door”] (b) Source: “Running-Water” $\rightarrow$ REQUIRED PLAN: [“nav”, “close_sink”] 3.
-OBJECT MAPPING INSTRUCTION (CRITICAL): You are provided with a partial hint about the audio content: One of the sound sources is CONFIRMED to be: ‘‘{obj_1}".
-You have to interact with it first.
-Your task is to identify the other sound source from the audio and generate the plan.
-You MUST map the objects exactly as follows: 1.
-Key “first_sound”: Generate the plan for ‘‘{obj_1}".
-2.
-Key “second_sound”: Identify the SECOND sound source from the audio (the one that is NOT ‘‘{obj_1}") and generate its plan.
-3.
-Do NOT output plans for any other potential objects unless they are detected as the second sound.
-4.
-CONSTRAINTS and OUTPUT FORMAT: 1.
-Output MUST be a valid JSON object containing a single root key “plan”.
-2.
-The keys inside “plan” MUST be exactly “first_sound” and “second_sound”.
-3.
-Do not include markdown formatting (like ‘‘‘ json), just the raw JSON string.
-5.
-Expected JSON Structure: { “plan”: { “first_sound”: [“skill_1”, “skill_2”, …], # $\leftarrow$ Actions for the object assigned to ’first_sound’ “second_sound”: [“skill_1”, “skill_2”, …] # $\leftarrow$ Actions for the object assigned to ’second_sound’ } }
-User Prompt: 1. Image Input 2. Audio Input 3. “Input observation provided. Based on the audio class and visual context, output the execution plan JSON.”
+Notably, in challenging dual-source scenarios, the baseline shows the capability to interact with the primary source amidst acoustic interference from a secondary signal, and successfully proceed to manipulate the subsequent target.